# See What I Mean? Expressiveness and Clarity in Robot Display Design

### v1

@@ I Introduction
Effective communication is critical in human-robot interaction (HRI), particularly in collaborative tasks where humans and robots must work together to achieve shared goals. For robots to communicate their intentions effectively, their actions and signals must be readable, predictable, and intuitive to humans [27]. In this context, visual displays play a vital role as they serve as the primary interface for robots to convey their internal states, intentions, and guidance in real-time [23]. When robots express intent clearly, they reduce the cognitive load on human collaborators, enabling faster and more accurate decision-making during tasks [3].
@@
Animation has emerged as a powerful tool in robotics for enhancing communication [23]. Animated displays add dynamism and temporal cues to robot communication, making their actions and signals more expressive and engaging. Animation techniques, such as blending, timing, and motion fluidity [1], can create a sense of purposeful movement that humans naturally interpret as indicative of internal states or goals. As a result, animation improves the readability and legibility of a robot’s actions, particularly in scenarios requiring quick and intuitive human responses [19].
Anthropomorphism in animation – the application of human-like characteristics to robots such as expressive eyes or gaze direction– is another widely explored means of enhancing communication in HRI [8]. Anthropomorphic features tend to evoke natural human responses, foster trust, relatability and perceived competence, and improve the user’s ability to interpret robot actions by mimicking familiar social cues such as attention, emotion, and intent [15].
@@ II Background
Non-verbal expressions have been widely used in human-robot communication [14][4][13][12], and have been shown to be successful in many applications, such as successfully influencing task performance and perception of robots [14]. Human participants improved their task performance when robots used non-verbal cues [27], even when they did not explicitly recognize or report the cues [14]. Our study is centered on the following designs in robotics: Anthropomorphism [16][11][21], Animation [19][23] and Non-verbal Communication in Collaborative Tasks [4][13][14].
@@ II-A Anthropomorphism in Robotics
Anthropomorphism in robotics refers to the design practice of giving robots human-like characteristics, such as facial features or gestures, to make them more relatable, intuitive and emotionally engaging [5][22]. This includes ensuring that robot outward expressions, such as facial movements [12], gaze shifts [8], and gestures [10], align with how humans naturally interpret social signals [4][5]. Humans naturally attribute human characteristics to non-human objects, including robots, which can make them more appealing and engaging [12]. Anthropomorphic features in product design can attract people, as even minimal social cues are enough to trigger anthropomorphism [17]. The design of the eyes and face of a robot significantly influences how people perceive and interact with it. Robots with more lifelike eyes are perceived as more personable and more suitable for the home environment [17]. The results of one notable study showed that robots with lifelike eyes were rated higher on personable traits (e.g., sensible, loyal, thoughtful) and were deemed more suitable for home environments, while abstract or no eyes were associated with professional traits but were perceived as cold and harder to interact with [16]. Not all studies, however, are in agreement that the addition of eyes improves human-robot communication and collaboration [18].
@@ II-B Animation in Robotics
Animation is an important advancement in robotic design, enhancing user engagement and social relatability [20]. Researchers use animation to try to make robot behaviors more readily apparent to human users [25] [4]. The integration of motion-based expressive elements enables robots to convey intent dynamically, often eliciting stronger human responses compared to static imagery. Past studies has described robotic animation as a mechanism for simulating life, wherein robots convey internal states, motivations, and responsiveness to maintain interactive autonomy [19]. Some researchers have concluded that animation enhances the perception of movement and intent, making emotionally expressive eye gestures easier to process quickly (The 12 Principles of Animation) [10]. The introduction of continuous, dynamic signals could increase cognitive processing demands, particularly in time-sensitive decision-making scenarios where users must extract clear, actionable information [23]. From an operational standpoint, it is critical to assess whether the illusion of life, as conveyed through animation, enhances or impairs the robot’s ability to provide precise navigational cues.
@@ II-C Non-Verbal Communication in Collaborative Tasks
Previous work in this field has highlighted the importance of readability and social understanding in the design of social robots [4][5]. To foster seamless interaction, robots must achieve real-time responsiveness, coherence in their movements, synchrony between actions and expressions, and expressive versatility to reflect emotions [26]. By making these social cues readable, robots can support humans’ social ability to understand others, thereby improving the quality and naturalness of interactions [24]. Building on the importance of anthropomorphism in fostering social interaction, studies examined how non-verbal cues, particularly gaze, could help humans accurately identify a robot’s intentions [8]. Semioticians have classified signs into three categories [9]: Symbol, which is an arbitrary signifier that is learned such as road signs; Icon, which is a signifier that resembles or mimics the signified such as a cartoon; and Index, which is a signifier that is not arbitrary such as a footprint. They can include a wide range of human-robot interaction cues, such as directional arrows or familiar traffic signals like red and green lights. Semiotic signs and symbols are useful as a framework to model human-robot interactions [7].
@@ III-D Procedure
Following the completion of both maze trials, participants filled out a post-task questionnaire. The questionnaire employed an 11-point Likert scale to measure trust [2], satisfaction [2], and understandability [6]. After completing the post-trial questionnaire, participants were asked to complete a second task we informally called the “Interpretation Game.” It consisted of a list of display signals corresponding to their assigned group of either Animated Displays or Static Displays. They were asked to interpret each signal and provide their best understanding of what the robot had tried to communicate via each display during the study. The evaluation of the Interpretation Game was conducted by the author and two other assistants. Each response was evaluated on an accuracy scale ranging from 0 to 2, where 0 indicated an incorrect interpretation, 1 represented a neutral or ambiguous response, and 2 signified a completely accurate interpretation. Participants who did not recall seeing a particular signal during the study were instructed to write, “I didn’t see the icon.” which did not factor into our calculation. Finally, the different graders’ scores for each participant’s responses were averaged for each of the responses to each display.
@@ IV-A Participants with static eyes modality had the highest percentage of task successes to task failures
[…]

### v2

@@ I Introduction
Effective communication is critical in human-robot interaction (HRI), particularly in collaborative tasks where humans and robots must work together to achieve shared goals. For robots to communicate their intentions effectively, their actions and signals must be readable, predictable, and intuitive to humans [s21175722]. In this context, visual displays play a vital role as they serve as the primary interface for robots to convey their internal states, intentions, and guidance in real-time [10.1145/3317325][fang2025demonstration]. When robots express intent clearly, they reduce the cognitive load on human collaborators, enabling faster and more accurate decision-making during tasks [10.3389/frobt.2022.754955].
@@
Animation has emerged as a powerful tool in robotics for enhancing communication [10.1145/3317325]. Animated displays add dynamism and temporal cues to robot communication, making their actions and signals more expressive and engaging. Animation techniques, such as blending, timing, and motion fluidity [10.7230/KOSCAS.2017.49.677], can create a sense of purposeful movement that humans naturally interpret as indicative of internal states or goals. As a result, animation improves the readability and legibility of a robot’s actions, particularly in scenarios requiring quick and intuitive human responses [10.1145/2157689.2157814].
Anthropomorphism in animation – the application of human-like characteristics to robots such as expressive eyes or gaze direction– is another widely explored means of enhancing communication in HRI [10.1145/3543174.3546841]. Anthropomorphic features tend to evoke natural human responses, foster trust, relatability and perceived competence, and improve the user’s ability to interpret robot actions by mimicking familiar social cues such as attention, emotion, and intent [10.1145/3319502.3374839].
@@ II Background
Non-verbal expressions have been widely used in human-robot communication [10.1145/1514095.1514110][1545011][10.1145/1514095.1514109][8525767], and have been shown to be successful in many applications, such as successfully influencing task performance and perception of robots [10.1145/1514095.1514110][yu2024much]. Human participants improved their task performance when robots used non-verbal cues [s21175722], even when they did not explicitly recognize or report the cues [10.1145/1514095.1514110]. Our study is centered on the following designs in robotics: Anthropomorphism [10.3389/frobt.2022.848295][10.1145/3171221.3171286][ROESLER2024108008], Animation [10.1145/2157689.2157814][10.1145/3317325] and Non-verbal Communication in Collaborative Tasks [1545011][10.1145/1514095.1514109][10.1145/1514095.1514110].
@@ II-A Anthropomorphism in Robotics
Anthropomorphism in robotics refers to the design practice of giving robots human-like characteristics, such as facial features or gestures, to make them more relatable, intuitive and emotionally engaging [10.5555/515422][saldien2014motion]. This includes ensuring that robot outward expressions, such as facial movements [8525767], gaze shifts [10.1145/3543174.3546841], and gestures [9900752], align with how humans naturally interpret social signals [1545011][10.5555/515422][yu2021active][tan2020top]. Humans naturally attribute human characteristics to non-human objects, including robots, which can make them more appealing and engaging [8525767]. Anthropomorphic features in product design can attract people, as even minimal social cues are enough to trigger anthropomorphism [10.3389/frobt.2023.1178433]. The design of the eyes and face of a robot significantly influences how people perceive and interact with it. Robots with more lifelike eyes are perceived as more personable and more suitable for the home environment [10.3389/frobt.2023.1178433]. The results of one notable study showed that robots with lifelike eyes were rated higher on personable traits (e.g., sensible, loyal, thoughtful) and were deemed more suitable for home environments, while abstract or no eyes were associated with professional traits but were perceived as cold and harder to interact with [10.3389/frobt.2022.848295]. Not all studies, however, are in agreement that the addition of eyes improves human-robot communication and collaboration [PILACINSKI2023e18164].
@@ II-B Animation in Robotics
Animation is an important advancement in robotic design, enhancing user engagement and social relatability [Ribeiro2020ThePO]. Researchers use animation to try to make robot behaviors more readily apparent to human users [6281390] [1545011]. The integration of motion-based expressive elements enables robots to convey intent dynamically, often eliciting stronger human responses compared to static imagery. Past studies has described robotic animation as a mechanism for simulating life, wherein robots convey internal states, motivations, and responsiveness to maintain interactive autonomy [10.1145/2157689.2157814]. Some researchers have concluded that animation enhances the perception of movement and intent, making emotionally expressive eye gestures easier to process quickly (The 12 Principles of Animation) [9900752]. The introduction of continuous, dynamic signals could increase cognitive processing demands, particularly in time-sensitive decision-making scenarios where users must extract clear, actionable information [10.1145/3317325]. From an operational standpoint, it is critical to assess whether the illusion of life, as conveyed through animation, enhances or impairs the robot’s ability to provide precise navigational cues.
@@ II-C Non-Verbal Communication in Collaborative Tasks
Previous work in this field has highlighted the importance of readability and social understanding in the design of social robots [1545011][10.5555/515422][fang2025charmconsideringhumanattributes][yu2023thumbs]. To foster seamless interaction, robots must achieve real-time responsiveness, coherence in their movements, synchrony between actions and expressions, and expressive versatility to reflect emotions [Breemen2004BringingRT]. By making these social cues readable, robots can support humans’ social ability to understand others, thereby improving the quality and naturalness of interactions [5650128]. Building on the importance of anthropomorphism in fostering social interaction, studies examined how non-verbal cues, particularly gaze, could help humans accurately identify a robot’s intentions [10.1145/3543174.3546841]. Semioticians have classified signs into three categories [crow2010visible]: Symbol, which is an arbitrary signifier that is learned such as road signs; Icon, which is a signifier that resembles or mimics the signified such as a cartoon; and Index, which is a signifier that is not arbitrary such as a footprint. They can include a wide range of human-robot interaction cues, such as directional arrows or familiar traffic signals like red and green lights. Semiotic signs and symbols are useful as a framework to model human-robot interactions [10.1007/978-3-031-38454-7_10].
@@ III-D Procedure
Following the completion of both maze trials, participants filled out a post-task questionnaire. The questionnaire employed an 11-point Likert scale to measure trust [bagheri2022transparent], satisfaction [bagheri2022transparent], and understandability [inbook]. After completing the post-trial questionnaire, participants were asked to complete a second task we informally called the “Interpretation Game.” It consisted of a list of display signals corresponding to their assigned group of either Animated Displays or Static Displays. They were asked to interpret each signal and provide their best understanding of what the robot had tried to communicate via each display during the study. The evaluation of the Interpretation Game was conducted by the author and two other assistants. Each response was evaluated on an accuracy scale ranging from 0 to 2, where 0 indicated an incorrect interpretation, 1 represented a neutral or ambiguous response, and 2 signified a completely accurate interpretation. Participants who did not recall seeing a particular signal during the study were instructed to write, “I didn’t see the icon.” which did not factor into our calculation. Finally, the different graders’ scores for each participant’s responses were averaged for each of the responses to each display.
@@ IV-A Participants with static eyes modality had the highest percentage of task successes to task failures
We calculated the mean scores per modality for their Interpretation Game: Animated Eyes 0.888, Animated Icons 1.321, Static Eyes 0.839 and Static Icons at 1.403. We found that static icons are most interpretable, using a one-way ANOVA which confirmed a significant difference across modalities (F(3,N4)=9.06,p<.001)(F(3,N-4)=9.06,p<.001). To evaluate whether this objective performance was reflected in participants’ self-perceived understanding, we computed Spearman rank correlations between the average Interpretation Game scores and the raw (-3 to 3) understandability ratings for each condition. Across all four modalities, we found no statistically significant correlation between participants’ self-assessments and their objective interpretation scores: Animated Eyes ($r_{s}$ = 0.281, p=0.1935), Animated Icons ($r_{s}$ =0.158, p = 0.4723), Static Eyes ($r_{s}$ = 0.487, p = 0.0772) and Static Icons ($r_{s}$ = 0.235, p = 0.4189)